Given the function H(z) = 8z+8 find each of the following:
H(7)=
H(-5)=
H(6.9)=

Answer by jim_thompson5910(35256)

H(z) = 8z+8


H(7) = 8(7)+8


H(7) = 56+8


H(7) = 64

-------------------------------------------------------
H(z) = 8z+8


H(-5) = 8(-5)+8


H(-5) = -40+8


H(-5) = -32


-------------------------------------------------------
H(z) = 8z+8


H(6.9) = 8(6.9)+8


H(6.9) = 55.2+8


H(6.9) = 63.2
-------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim

-------------------------------------------------------------------------------------------------